OpenStudy (thesecret20111):

Can someone help with this? 1/cos^2a+1-3tan^2a=0

OpenStudy (mertsj):

\[\frac{1}{\cos ^2a}+1-3\tan ^2a=0\]

OpenStudy (mertsj):

Maybe that's it.

OpenStudy (thesecret20111):

Yes your second one is the question I'm given.

OpenStudy (mertsj):

\[\frac{1}{\cos ^2a}+\frac{\cos ^2a}{\cos ^2a}-\frac{3\sin ^2a}{\cos ^2a}=0\]

OpenStudy (mertsj):

Does that help?

OpenStudy (mertsj):

I would now add the fractions and replace sin^2 with 1-cos^2

OpenStudy (thesecret20111):

How did you get the middle part?

OpenStudy (thesecret20111):

Oh from +1

OpenStudy (mertsj):

yes

OpenStudy (thesecret20111):

so 1cos^2a/cos^2a-3sin^2a/cos^2a=0

OpenStudy (thesecret20111):

Then 1cos^2a/cos^2a- 3(1-cos^2)/cos^2a=0?

OpenStudy (thesecret20111):

Ok I think I can get it form here.. Thanks so much for the help!

OpenStudy (mertsj):

yw

OpenStudy (mertsj):

Make sure that is 1+cos^2a not 1cos^2a

OpenStudy (mertsj):

\[\frac{1+\cos ^2a-3(1-\cos ^2a)}{\cos ^2a}=0\]

OpenStudy (mertsj):

\[\frac{1+\cos ^2a-3+3\cos ^2a}{\cos ^2a}=0\]

OpenStudy (mertsj):

\[\frac{4\cos ^2a-2}{\cos ^2a}=0\]

OpenStudy (mertsj):

So you are saying that: \[\frac{3(1-\cos ^2a)}{\cos ^2a}=0\]

OpenStudy (mertsj):

That does not equal 0

OpenStudy (mertsj):

\[3\tan ^2a=3\frac{\sin ^2a}{\cos ^2a}=\frac{3}{1}\frac{\sin ^2a}{\cos ^2a}=\frac{3(1-\cos ^2a)}{\cos ^2a}\]

OpenStudy (mertsj):

They are the same thing.

OpenStudy (mertsj):

no

OpenStudy (mertsj):

\[\frac{1-\cos ^2a}{\cos ^2a}=\frac{1}{\cos ^2a}-\frac{\cos ^2a}{\cos ^2a}\]

OpenStudy (mertsj):

You lost me. I thought we had a equation.

OpenStudy (thesecret20111):

Yeah I think I need to go back a lot of steps, I must have made a mistake somewhere. :(

OpenStudy (thesecret20111):

1+cos^2a/cos^2a-3(1-cos^2)/cos^2a=0\[\frac{ 1+\cos^2a }{ \cos^2a } -\frac{ 3(1-\cos^2a) }{ \cos^2a }=0\] I think thats what I'm up to.

OpenStudy (mertsj):

Yes. I agree.

OpenStudy (thesecret20111):

So do I add these fractions?

OpenStudy (mertsj):

That's what I did. Or I guess you could put the one fraction on the right side and then cross multiply.

OpenStudy (thesecret20111):

so would that be \[\frac{ 2(\sin a)^2+2 }{ \cos^2a }=0\]

OpenStudy (mertsj):

Where did the sin come from? We had all cos

OpenStudy (mertsj):

\[\frac{1+\cos ^2a-3(1-\cos ^2a)}{\cos ^2a}=0\]

OpenStudy (mertsj):

\[\frac{1+\cos ^2a-3+3\cos ^2a}{\cos ^2a}=0\]

OpenStudy (mertsj):

\[\frac{4\cos ^2a-2}{\cos ^2a}=0\]

OpenStudy (mertsj):

If a fraction is 0, its numerator is 0. Therefore: \[4\cos ^2a-2=0\]

OpenStudy (mertsj):

\[4\cos ^2a=2\] \[\cos ^2a=\frac{1}{2}\] \[\cos a=\pm\frac{\sqrt{2}}{2}\] \[a=\frac{\pi}{4}+\frac{\pi}{4}n\]

OpenStudy (thesecret20111):

Ok that makes sense. I obviously need a lot more trigonometry practise though :/ Thanks sooooo much for the help you're awesome!

OpenStudy (mertsj):

yw